Search
| Date | Author | Entry | File | Actions |
|---|---|---|---|---|
| 2025-11-21 | vanderws |
Sieve MLE WC - Hancock soil, sampled 11/19/2025 - Jon Filter LTER leachate - Alyssa Filter WN soil extracts - Olivia, Brian |
||
| 2025-11-21 | Alyssa |
Prepared Standards, Lachat run LTAR Lysimeter 20250624. |
||
| 2025-11-21 | vanderws |
Collect leaf litter from traps in T5, G8, CF, DF, SF. Pull traps from G8 as trees are bare. - Jon, Brian, Olivia Pull T6 station flags - Jon |
||
| 2025-11-20 | vanderws |
Sieve MLE WN - Rhinelander soil, sampled 11/18/2025. Weigh dry soil tins - Stacey Weigh WN soil for extraction and moisture - Stacey, Alyssa |
||
| 2025-11-20 | sven |
Adjusted the charge parameters on tower 1 to the default parameters for Sealed Lead Acid batteries. I added a battery to power the access point prior to programming the charge controller. While that allowed me to finish programing the controller, it still required a power cycle to restore power to the load. |
||
| 2025-11-20 | Alyssa |
Prepared 1M KCL. Finished dilutions from LTAR Lysimeter 20250610. |
||
| 2025-11-20 | vanderws |
BCSE Series 22 gas sampling - Jon, Brian, Alyssa (Hydrosense) Collect LTER leachate - Brian, Jon |
||
| 2025-11-19 | vanderws |
Evacuate LTER lysimeters - Jon, Olivia BCSE gas chamber prep - Brian |
||
| 2025-11-19 | kahmark |
ASP tower 1 rain gauge tipping mechanism was not working. Spider web was preventing it from tipping. Repaired. Missed yesterdays light rain. Added camera to tower 2. made final adjustment of height (2m) and direction (N-S) on each tower KK DW |
||
| 2025-11-19 | vanderws |
BCSE gas vial, tote prep - Brian Filter LTER soil extracts - Jon, Brian |
||
| 2025-11-19 | Alyssa |
Lachat run LTAR Lysimeter 20250610. |
||
| 2025-11-18 | Alyssa |
Prepared 1M KCL, Lachat run LTAR Lysimeter 20250529. Prepared Color Reagent (2 L) 200 ml phosphoric acid 80 gm sulfanilamide, 2.0 gm N-1-naphthylethylenediamine dihydrochloride, Dilute to 2L with Nanopure H2O. |
||
| 2025-11-18 | vanderws |
Weigh LTER soil for moisture and nitrogen extraction - Brian, Jon Pack CN - Brian Dishes - Jon |
||
| 2025-11-17 | vanderws |
LTER soil sampling - Alyssa, Brian, Olivia, Jon |
||
| 2025-11-17 | dave |
GC Sequence LTAR20251113S28BGC had a few vials submitted out of order. All samples run sequentially on GC. Vials submitted out of order as follows: …62, 63, 68, 65, 66, 67, 64, 69, 70… Also, note on tape says vials 41 and 44 were swapped, though they were submitted in sequential order in racks. |
||
| 2025-11-17 | vanderws |
Sieve LTER soil - Alyssa, Brian, Olivia, Jon |
||
| 2025-11-14 | Alyssa |
Reran Channel 2 and dilutions of Lachat run LTAR Lysimeter 20250513. |
||
| 2025-11-14 | vanderws |
Sieved BCSE soil and set out to air-dry - Sarah Sieved Sorghum Nitrogen Rate soil and subsampled to send to A&L for S1 analysis. - Slater, Jon Filter LTAR leachate, dishes - |
||
| 2025-11-14 | vanderws |
Replace BCSE then LTAR station flags - Olivia, Brian Sample Sorghum Nitrogen Rate Experiment soil. Location 2: one core per plot, pooled by block, 4 total. Location 1: split into northern and southern halves, one core per plot, pooled by block, 8 total. - Slater, Jon Sample BCSE soil for laboratory respiration incubation. G1,3,5,8,10, R1-3. Four cores/plot, one or two per station, alternating within and between crop row for G1, G3 and G8. Three cores/plot, one per station pooled for G5 and G10. - Sarah |
||
| 2025-11-13 | vanderws |
Measure LTAR leachate - Slater, Brian |
||
| 2025-11-13 | vanderws |
LTAR Series 28 gas sampling - Slater, Brian, Jon, Alyssa (Hydrosense) Collect LTAR leachate - Slater, Brian |
||
| 2025-11-13 | Alyssa |
Replaced Cadmium Column and reconditioned. Lachat run LTAR Lysimeter 20250513. Notes from run:
Issue with calibration curve/column in channel 2 results - rerun 11/14. |
||
| 2025-11-12 | vanderws |
LTAR gas chamber prep - Olivia, Sarah Evacuate LTAR lysimeters; replace BCSE station flags - Slater, Jon |
||
| 2025-11-12 | vanderws |
LTAR gas vial, tote prep - Olivia, Sarah Weigh dry tins, dishes - Sarah |
||
| 2025-11-12 | Alyssa |
Acid washed dishes, Prepared standards. Lachat run LTAR Lysimeter 20250428. Notes from run:
|
||
| 2025-11-11 | Alyssa |
Prepared Color Reagent (2 L) 200 ml phosphoric acid 80 gm sulfanilamide, 2.0 gm N-1-naphthylethylenediamine dihydrochloride, Dilute to 2L with Nanopure H2O. Prepared Ammonium Chloride Reagent (4L): 320 gms Ammonium Chloride 4.0 gms EDTA pH to 8.45 with NaOH, store Lachat run LTAR Lysimeter 20250415. |
||
| 2025-11-11 | vanderws |
Put away clean, dry dishes, weigh remaining BCSE grain, grind SWF 2024 harvest plant samples - Jon |
||
| 2025-11-10 | vanderws |
LTER Series 15 gas sampling - Alyssa, Slater, Jon, Olivia, Sarah LTER Hydrosense pm - Alyssa |
||
| 2025-11-10 | vanderws |
Finish sorting ASP2 plants - Sarah |
||
| 2025-11-10 | sven |
Attempted to use the MSView application to increase the Target voltage on tower 1 from 14.2 to 14.6 and the float voltage from 13.2 to 13.7. I had lowered these values for the LI-Ion battery, but now that we have an AGM battery they should probably be set back. Unfortunately the Sunsaver cuts power when writing to the coils, so the network goes down after the first coil is written too. I rebooted the Sunsaver and the Target voltage was set but the float voltage was still at the old value. I the future, I will need to provide an alternate power source for the tower router before attempting to set the parameters through the MSView application. |
||
| 2025-11-07 | Alyssa |
Finished dilutions from Lachat run LTAR Lysimeter 20250401. |
||
| 2025-11-07 | vanderws |
Collect leaf litter from traps in T5, G8, CF, DF, SF; reinstall LTER gas chambers in T1-4, plus prairie strips - Brian, Slater, Jon, Olivia Shell LTAR Inhibitor Study corn with Josh Dykstra - Sarah |
||
| 2025-11-06 | Alyssa |
Finished dilutions from Lachat run LTAR Lysimeter 20250318. Acid washed dishes. Prepared fresh Phosphate Buffer (1L) 28 gms Sodium Hydroxide pellets 50 gms sodium phosphate dibasic heptahydrate dilute to 1 L with nanopure. Prepared EDTA Reagent (1L) 66 gms ethylenediamine tetraacetic acid disodium salt dihydrate (EDTA) Approximately 7.0 gms sodium hydroxide pellets ph to 7.03 with NaOH pellets then dilute to volume with nanopure. Lachat run LTAR Lysimeter 20250401. Notes from run:
|
||
| 2025-11-06 | vanderws |
Sort ASP2 plants, put LTAR 11/3/2025 soil out to air-dry - Slater, Brian, Jon |
||
| 2025-11-05 | vanderws |
Sort ASP2 plants - Slater, Brian, Olivia Finish BAF subsampling for 2024 BCSE peak plant CN - Sarah Dishes - Jon Filter LTAR extracts, dishes - Slater, Brian, Jon |
||
| 2025-11-05 | vanderws |
Sample LTAR Inhibitor Study corn - Kevin, Dave, Sarah, Josh |
||
| 2025-11-05 | Alyssa |
Finished dilutions from Lachat run LTAR Lysimeter 20250303. Lachat run LTAR Lysimeter 20250318. Notes from run:
Reconditioned Column. |
||
| 2025-11-04 | kahmark |
BAU towers raised for tillage at 12:30 today. No sensors yet on tower 4. Tower 2 CS650 sensor 4 chewed in half again. |
||
| 2025-11-04 | Alyssa |
Prepared 1M KCL. |
||
| 2025-11-04 | vanderws |
Prep LTER gas chambers. After chambers were prepped, heard from Joe that gas sampling could not happen tomorrow due to lime applications planned for T1-4. - Slater, Brian Pulled all T1-4, plus prairie strip, gas chambers for lime. Finished clipping ASP2 plants. - Alyssa, Slater, Brian, Jon |
||
| 2025-11-04 | vanderws |
Weigh LTAR soil for extraction and moisture - Alyssa, Jon Prep LTER gas vials, totes - Alyssa, Jon, Brian Sort ASP2 plants - Alyssa, Slater, Brian, Jon |
||
| 2025-11-03 | Alyssa |
Prepared Standards. Prepared Color Reagent (2 L) 200 ml phosphoric acid 80 gm sulfanilamide, 2.0 gm N-1-naphthylethylenediamine dihydrochloride, Dilute to 2L with Nanopure H2O. Lachat run LTAR Lysimeter 20250303. Notes from run:
|
||
| 2025-11-03 | vanderws |
Sieve and subsample LTAR soil for A&L - Jon, Brian, Olivia, Sarah Weigh dry soil tins, dishes - Slater |
||
| 2025-11-03 | vanderws |
LTAR November soil sampling, 4 cores/station, 12 cores total to ensure enough soil for both A&L analysis and air-dried archive besides extraction and moisture - Slater, Jon, Brian, Olivia, Sarah Pull ASP6 flags and chambers for harvest Start ASP2 plant sampling; Stations 1, 2, 3: 2mS, 0.5mW to NE corner of 2 x 0.5m frame, oriented with the 0.5m ends on the east and west. R5: Split the plot into quarters. Place the frame approximately in the center of the western most quarter (label west) and the eastern most quarter (label east). The frame should be perpendicular to forage rows, if visible. Sorted later in the field lab to MEDSA, TRFPR, CICSP, DACGL, BRSNA (live volunteer canola), and remainder UNSRT. - Jon, Brian, Olivia, Sarah |
||
| 2025-10-31 | dave |
BCSE20251023S21BGC sequence notes: All samples run in sequential vial number order. Samples submitted out of order in trays as follows: …59, 60, 61, 78, 79, 64, 65, 74, 75, 68, 69, 70, 71, 72, 73, 66, 67, 76, 77, 62, 63, 80… |
||
| 2025-10-31 | Alyssa |
Lachat run LTAR Soil 20250512. |
||
| 2025-10-31 | vanderws |
Sieve SWF soil and put a portion in A&L soil bags to send for agronomic analysis. - Slater, Jon, Brian, Emily Sort ASP1 plants - Ainsley, Olivia |
||
| 2025-10-31 | vanderws |
Switchgrass Nitrogen Rate Experiment (SWF) soil sampling to send to A&L. Six cores/plot taken from the yield strip harvested area down the center of each plot, along a diagonal from NE to SW plot corners. - Brian, Slater, Jon, Emily |
||
| 2025-10-31 | kahmark |
Lowered tower 1 to 2 m above the soil. Cleaned/unplugged rain gauge at 11:20ish. Reinstalled all ground sensors. |
||
| 2025-10-30 | vanderws |
Weigh plants - Brian Grind plants - Slater Sort ASP1 plants - Alyssa, Ainsley, Brian, Emily, Jon, Slater |
||
| 2025-10-30 | dave |
Fix on LTAR lysimeter bottles not yielding/not holding vacuum. Site BAU2, Rep 2 bottles checked - lines between bottle and prenart appear to be intact, line connection loose at bottle attachment fixed/sealed, o-rings reversed to proper positioning. Fixed O-rings and one pinched tube on ASP5, Rep 2. Lines appear to be intact between bottles and prenarts. |
||
| 2025-10-30 | vanderws |
LTAR 27 gas sampling - Emily, Ainsley, Jon, Alyssa (Hydrosense) ASP1 plant sampling ahead of mowing tomorrow, Stations 1, 2, 3: 1mS, 0.5mW to NE corner of 2 x 0.5m frame, oriented with the 0.5m ends on the east and west. R5: Place the frame approximately in the center of each half of the plot (east and west halves). Sorted in the field lab later… MEDSA, TRFPR, CICSP, DACGL, BRSNA (live volunteer canola), remainder UNSRT - Alyssa, Slater, Brian, Jon, Emily, Ainsley |
||
| 2025-10-29 | Alyssa |
Lachat run MDARD August 2025 Soil. |
||
| 2025-10-29 | kahmark |
Sensors back in BAU1. Replaced soil moisture sensors and 3 heat flux plates. We need more 107s or 616s for soil temp. |
||
| 2025-10-29 | vanderws |
LTAR gas vial, tote prep - Ainsley, Slater Weigh plants - Brian Filter MLE MN soil extracts - Ainsley, Jon |
||
| 2025-10-29 | vanderws |
LTAR gas chamber prep - Ainsley, Slater T7 post frost plant sampling, 2mS, 4mW of each station flag to the NE corner of 2x0.5m frame - Olivia, Sarah, Emily, Jon, Slater, Brian |
||
| 2025-10-28 | vanderws |
Weigh MLE MN soil for extraction and moisture - Ainsley, Brian Measure and filter LTER leachate - Alyssa, Emily Pack CN - Ainsley BAF subsampling for BCSE 2024 - Jon Grind plants - Slater Weigh plants - Brian |
||
| 2025-10-28 | sven |
updated flux tower 2 program to reflect the new heat flux plate calibrations. The calibration coefficients for heat flux plate 1 and 3 needed to be reversed to match the way the probes were hooked up. |
||
| 2025-10-28 | vanderws |
Collect LTER leachate - Alyssa, Slater Help Levi Dreiling with barley sampling near the Rainout Building, field D2 - Emily, Jon |
||
| 2025-10-27 | vanderws |
Evacuate LTER lysimeters - Jon, Emily Pull G4, 5, 7, 9, 10 flags and frames for harvest - Olivia, Sarah Help Mark move OTCs - Emily, Jon |
||
| 2025-10-27 | Alyssa |
Acid washed dishes, Lachat run LTAR Soil 20250414. |
||
| 2025-10-27 | vanderws |
Dishes, weigh dry soil tins - Slater, Ainsley Weigh plants - Sarah Grind plants - Jon Sieve MLE MN soil (sampled 10/23/2025) - Ainsley, Slater, Emily |
||
| 2025-10-27 | kahmark |
Removed tower 1 ground sensors for harvest at 10:15AM. Looks like the rain gauge is completely full of water but too high (14.5’) to really tell.. |
||
| 2025-10-24 | Alyssa |
Prepared Standards, Lachat run LTAR Soil 2025031. |
||
| 2025-10-24 | vanderws |
Collect leaf litter from T5, G8, CF, DF, SF - Slater, Emily, Jon |
||
| 2025-10-24 | vanderws |
Weigh plants - Jon, Ainsley Pack CN - Ainsley, Slater |
||
| 2025-10-23 | Alyssa |
Replaced all Lachat reagent tubing. |
||
| 2025-10-23 | vanderws |
BCSE 21 gas sampling - Emily, Jon, Alyssa (Hydrosense) |
||
| 2025-10-23 | vanderws |
Weigh plants - Slater Pack CN - Ainsley, Jon Filter MLE MS soil extracts - Emily, Ainsley |
||
| 2025-10-22 | dave |
LTAR20251016S26BGC GC sequence run had sample vials submitted out of order. All sample vials run on GC in numerical order. Submitted vials out of order as follows: …96, 97, 118, 119, 120, 101, 122, 123-144, 117, 98, 99, 100, 121, 102, 103-116. |
||
| 2025-10-22 | Alyssa |
Prepared 1M KCL |
||
| 2025-10-22 | vanderws |
Collect LTAR leachate - Slater, Jon BCSE gas chamber prep - Sarah |
||
| 2025-10-22 | dave |
LTER20251010S14BGC samples submitted out of sequence. All sample vials run in numerical order, but they were submitted in this order: …161, 162, 167, 164, 165, 166, 163, 168, 169, 170… then again, …219, 220, 222, 221, 223, 224, 225… |
||
| 2025-10-22 | vanderws |
Measure and filter LTAR leachate - Slater, Jon BCSE gas vial, tote prep - Sarah Weigh MLE MS soil for extraction and moisture - Ainsley, Emily Weigh plants - Olivia, Ainsley Pack CN - Slater Data entry - |
||
| 2025-10-21 | vanderws |
Evacuate LTAR lysimeters - Ainsley, Jon Shell and shred G1 corn with Jarrod/Ethan operating the shredder - Slater, Emily |
||
| 2025-10-21 | vanderws |
Weigh plants - Alyssa, Ainsley Pack CN - Sarah, Alyssa Sieve MLE MS soil (sampled 10/21 am by Todd and Chris) - Slater, Emily, Jon |
||
| 2025-10-20 | vanderws |
Pack CN - Olivia, Jon Weigh plants - Slater, Ainsley Dishes - Slater ChemInventory spreadsheet set up - Alyssa |
||
| 2025-10-20 | vanderws |
Shell and shred LTAR corn with Josh Dykstra operating the shredder - Jon, Ainsley Shell IP1 corn, no shredding/stover discarded per Brook’s instructions - Emily, Sarah Dump soil - Sarah, Emily |
||
| 2025-10-17 | vanderws |
Finish checking transect data - Slater Weigh plants - Emily, Alyssa Pack CN - Alyssa, Brian Filter LTER soil extracts - Ainsley, Emily |
||
| 2025-10-17 | vanderws |
Finish N-deposition fertilization, SF2, DF3, and CF3 - Brian, Ainsley |
||
| 2025-10-17 | kahmark |
Wires trenched in at both new LTAR new towers. Waiting for sensors. |
||
| 2025-10-16 | vanderws |
LTAR Series 26 gas sampling. When ASP3 (including no compost microplot) and BAU2 chambers were installed on 10/15/2025, the corn stover was removed for installation and not replaced. This was noted during sampling but stover wasn’t replaced inside those chambers until after sampling today. - Jon, Slater, Emily, Alyssa (Hydrosense) Fertilize N-deposition microplots, finished CF1, CF2, DF1, DF2, SF1, and SF3 - Brian, Emily, Olivia |
||
| 2025-10-16 | vanderws |
Weigh LTER soil for extraction and moisture - Olivia, Ainsley Weigh plants - Brian Transect data entry/checking - Ainsley, Alyssa Pack CN - Slater, Stacey |
||
| 2025-10-16 | Alyssa |
Uploaded GLBRC Scale up Lux Arbor and Marshall samples in multiple data sets due to partial contamination. Notes are as follows: L1 and L3 samples from the Lux Arbor Scale up field (12/2/2024) had Ammonia contamination in KCL used, so all ammonia results were discounted. L2 was sampled on different date (12/16/2024) and had no contamination, results will be uploaded in a separate file. M1, M2 and M4 samples from the Marshall Scale up field (12/2/2024) had Ammonia contamination in KCL used, so all ammonia results were discounted. M3 was sampled on different date (12/16/2024) and had no contamination, results will be uploaded in a separate file. |
||
| 2025-10-16 | Alyssa |
Prepared 1M KCL |
||
| 2025-10-16 | kahmark |
ASP Tower 2 and Bau tower 2 both infrastructures installed over the course of the week. Tower positions measured by Brook Wilke, with Kevin present. Sensors will come from Jiquan Chen. |
||
| 2025-10-15 | kahmark |
Dropped BAU Tower 1 to standard 2m height after corn harvest. Will take some time to move ground sensors back as 8/9 destroyed by coyotes. Tower height was very easy to drop down. |
||
| 2025-10-15 | vanderws |
LTER October soil sampling - Alyssa, Slater, Olivia, Brian, Ainsley, Jon LTAR gas chamber prep - Emily, Sarah |
||
| 2025-10-15 | vanderws |
LTAR gas vial, tote prep - Emily, Sarah Weigh plants - Ainsley Transect data entry/checking - Emily, Sarah, Alyssa, Olivia Pack CN - Sarah, Stacey Weigh dry soil tins, dishes - Brian Sieve LTER soil - Slater, Emily, Jon |
||
| 2025-10-14 | vanderws |
Transect data entry/checking - Brian, Ainsley, Emily, Slater Weigh G1 corn - Alyssa Dishes - Brian Pack CN - Stacey |
||
| 2025-10-14 | vanderws |
Thresh soybeans, canola - Emily, Slater, Alyssa, Ainsley Pull chambers, flags from ASP3 and BAU2 for corn harvest - Alyssa |
||
| 2025-10-13 | vanderws |
ASP3 corn sampling (same as BAU2): R1-4 all stations: 0.66mS, 0.66mW to NE corner of 1.52 x 0.66m frame, oriented with the 0.66m ends on the east and west. Perpendicular to corn rows. R5: Place the frame approximately in the center of each half of the plot (east and west halves), with the 0.66m frame ends oriented towards the north and south. - Slater, Brain, Olivia, Sarah, Emily, Ainsley Innovations Plot 1 (IP1) Short Corn Trial sampling: Area is split into East and West halves. Each half includes the same three treatments: Short Corn, Control 1 and Control 2. Six samples will be taken from each treatment, three from the east (labeled south, center, north) and three from the west (labeled south, center, north). The IP1 field is 85m from south to north, same as ACSE plots. So, the same distances from the south end were used to match the station flag distances from the south end: south was about 12m from the south end, center about 45 m from the south end, and north about 64m from the south end. Per Brook’s instructions on the EW sampling locations, short corn was sampled near the center of the eastern half of each “rep”, otherwise when planted in 3 rows - rows 2 and 3 from the west were sampled and when planted in 6 rows - rows 5 and 6 from the west were sampled. Use 1.52 x 0.66m frame, oriented with the 0.66m ends on the east and west. Perpendicular to corn rows.- Slater, Brain, Olivia, Sarah, Emily, Ainsley, Alyssa |
|
|
| 2025-10-13 | kahmark |
BAU Tower 79-2 ground sensors removed for harvest this morning. Found all sensors destroyed even after we covered them with corrugated cover. Coyote went right to the sensor and chewed off in all cases. |
||
| 2025-10-10 | vanderws |
Filter LTAR soil extracts, dishes - Slater, Emily Pack CN - Stacey |
||
| 2025-10-10 | vanderws |
LTER Series 14 gas sampling - Brian, Jon, Ainsley, Slater, Emily LTER Hydrosense - Alyssa |
||
| 2025-10-09 | vanderws |
Weigh LTAR soil for extraction and moisture - Brian, Ainsley LTER gas vial and tote prep - Slater Transect data entry - Jon Pack CN - Stacey |
||
| 2025-10-09 | vanderws |
LTER gas chamber prep - Collect leaf litter from traps in T5, G8, CF, DF, SF - |
||
| 2025-10-08 | vanderws |
LTAR October soil sampling - Thresh soybeans - |
||
| 2025-10-08 | vanderws |
Transect data entry - Brian, Ainsley Weigh plants - Sarah Pack CN - Stacey |
||
| 2025-10-07 | vanderws |
Pack CN - Stacey Weigh plants - Brian, Jon Plant species composition transect data entry - Emily, Slater, Ainsley, Brian Weigh dry soil tins, dishes - Jon |